Six years ago, Bradford Community Hospital issued 20-year municipal bonds with a 7% annual coupon rate. The bonds were called today for a $70 call premium—that is, bondholders received $1070 for each bond. What is the realized rate of return for those investors who bought the bonds for $1000 when they were issued?
Answer #1

Using financial calculator
N=6
FV=1070
PV=-1000
PMT=1000*7%=70
CPT I/Y=7.95528%
Realized return=7.955%